# Optimal Data Splitting for Holdout Cross-Validation in Large Covariance Matrix Estimation

### v1

@@ Abstract
Cross-validation is a statistical tool that can be used to improve large covariance matrix estimation. Although its efficiency is observed in practical applications, the theoretical reasons behind it remain largely intuitive, with formal proofs currently lacking. To carry on analytical analysis, we focus on the holdout method, a single iteration of cross-validation, rather than the traditional $k$-fold approach. We derive a closed-form expression for the estimation error when the population matrix follows a white inverse Wishart distribution, and we observe the optimal train-test split scales as the square root of the matrix dimension. For general population matrices, we connected the error to the variance of eigenvalues distribution, but approximations are necessary. Interestingly, in the high-dimensional asymptotic regime, both the holdout and $k$-fold cross-validation methods converge to the optimal estimator when the train-test ratio scales with the square root of the matrix dimension.
@@ 1 Introduction
Covariance estimation is a major question in statistics that finds applications in many areas such as physics [1], neuroscience [2][3], signal processing [4] and finance [5, 6, 7]. A simple answer could be to use the sample estimator; however, it is largely inefficient in high dimensions when the order of the matrix $n$ is comparable to the number of observations $t$ of the data [6]. More generally, it is known that the amount of information lost when the sample estimator $\bm{E}$ is used to approximate the population covariance $\bm{C}$ depends only on $q=\frac{n}{t}$ for any population matrix with Gaussian data in the high-dimension setting [8]. The estimator coincides with the population when $q=0$, then the larger is $q$ the larger is the error [6, 9].
@@
To address the lack of robustness in the sample covariance estimator, researchers have extensively studied Rotational Invariant Estimators (RIE) [10, 11, 12]. Today, RIEs represent the most widely used approach for covariance cleaning. In simple terms, a RIE is built by maintaining the eigenvectors of the sample covariance and correcting the sample eigenvalues. The reasoning is that when no information is available on the direction of the eigenvectors of the population covariance matrix, one can assume that each basis of eigenvectors has the same probability of appearing, therefore one keeps the sample eigenvectors and focuses on estimating the eigenvalues such that it minimizes a function of choice such as the Mean Square Error (MSE) [11], the Kullback-Leibler divergence [13] or the Stein’s loss [14]. In the last decades, many RIEs linked with Random Matrix Theory (RMT) were proposed to clean the noise of the sample estimator (see [10] and [9] for a review). One of the first RMT methods for cleaning is known as eigenvalue clipping. This method suggests that all the eigenvalues outside the sharp limit $(\lambda_{-},\lambda_{+})$ of the Marcenko-Pastur (MP) distribution should be kept, while the ones that fall within such limit should be regarded as noise and should be shrunk to a single average value [5, 15]. However, the limits of the MP distribution depend only on the measured aspect ratio $q$, i.e., the size of the matrix $n$ divided by the number of data points $t$. Notably, the measured aspect ratio is not always equal to the effective ratio that best fits the empirical spectral density to the MP distribution [5, 10]. The other widely used method for cleaning covariance matrices is linear shrinkage. The idea is to linearly shrink the sample eigenvalues towards the identity [6, 16]; although this approach works in many practical applications, it is theoretically justified only under some circumstances such as when the population covariance is a white inverse Wishart [16]. More recently, the optimal shrinkage correction was analytically derived by Ledoit and Péché under certain assumptions, including the $12$th absolute central moment finite for the noise and stationary data [11]; however, the computation of this quantity from their equation was not straightforward. Numerical solutions, known as Non-Linear Shrinkage (NLS), were implemented and further developed by Ledoit and Wolf [17, 18], making it more practical to apply. The NLS, under the high-dimensional limit, converges to the oracle estimator, which is the ideal estimator that minimizes the MSE with the population covariance matrix. It is worth remarking that the direct estimation of the oracle is unattainable in practice since it relies on the knowledge of the population covariance matrix, which is unknown [11].
@@
In practical scenarios, where the data may not be Gaussian or stationary, alternative methods inspired by machine learning find their best regime of applicability. Techniques such as Cross-Validation (CV), some based on factor models [19, 20], and others employing Bayesian methods with prior distributions for estimating the covariance matrix [21] can, in principle, outperform theoretically optimal methods, such as those derived by Ledoit and Péché, under these complex scenarios. However, there is no theoretical evidence to support this claim, as deriving analytical results for these methods is highly complex. To fix this gap, the focus of this work is specifically on covariance cleaning through CV.
@@
CV is one of the most widely used machine learning methods for model selection and evaluation and belongs to the family of Monte-Carlo methods (see [22] for a review of the main CV procedures). Its applications include a wide range of areas such as ecology [23], medical imaging [24], psychology [25], and finance [26], to quote a few. Part of the reason CV is widely used is because it almost needs no assumptions on the data, it helps reduce the risk of over-fitting for model selection and its efficiency is undoubtedly in practice [22, 26]. There are several CV procedures; however, the core principle of any of those involves dividing the dataset into separate training and testing subsets, also called in-sample and out-of-sample. The model is trained on the training subset and then evaluated on the testing subset. This process helps to assess how well the model generalizes to unseen data and is particularly useful to detect overfitting [27]. When overfitting occurs, the model may perform well on the training set but will show a noticeable drop in performance on the testing set. The most popular CV methods are the $k$-fold CV, the Leave-One-Out (LOO) CV, which is a particular case of the former [22], and the holdout, also called validation or simple cross-validation [22]. The holdout procedure consists of splitting a single time the data set into a train set and a test set. Therefore, this method consists of a single round of a $k$-fold CV. The recommendation in the literature is to use $10$ to $30\%$ of the data points for the test set and the rest for the train set [22]. Differently from the $k$-fold and LOO CV, the holdout can respect the causality of time series if the train and test data are ordered and if the test data is chosen posterior to the train. Practitioners should choose their approach based on the system specificity and the estimation strategy they aim for. For example, the LOO CV is adequate for bias correction but has a very high variance. On the other hand, if stability is important, one can run a $k$-fold procedure with a small k usually between $5$ and $10$, with $k=10$ being the most common one [28, 29, 30].
@@
A covariance cleaning approach inspired by CV was proposed in reference ([31]), and some authors [32, 33] have guessed that this estimator should converge to the optimal covariance cleaning method in the high-dimensional setting [32]. Other authors claim that the $k$-fold CV outperforms the NLS when applied to financial data [31, 34]. The flexibility of this numerical estimation allows for simple modifications such as including exponential smoothing in CV [26], or pre-filtering the data using hierarchical clustering [35]. In its original formulation, a LOO CV was proposed [17]; however, reference ([31]) reasoned that a $k$-fold would improve the estimation. The current procedure consists of dividing the dataset into $k$-folds and finding the average eigenvalue correction of the train eigenvalues to best fit the test set. Analogously to the $k$-fold problem, as $k$ increases, the train set becomes larger, but the test set shrinks, leading to a higher generalization variance. Finding the optimal way to split the data is challenging, especially because the underlying theory is not well understood. In the literature, the recommendation is again to use $k$ between $5$ and $10$ [31, 32, 26].
@@
In this paper, we present the first step toward gaining a better understanding of the performance of CV for large covariance matrix estimation using random matrix theory tools. We compute the element-wise MSE between the estimator and the population covariance of a single round for a $k$-fold CV, which coincides with the Frobenius error of the holdout estimator. For the general case, the error requires numerical approximations [17, 18]. To carry on this estimation analytically, instead, we assume a white inverse Wishart distributed population covariances. When the population covariance is an inverse Wishart, both the Bayesian estimator and linear shrinkage coincide with the oracle [36, 37]. We believe that studying the holdout method is a necessary step to understand the $k$-fold CV better. Furthermore, while the holdout method may be less efficient than $k$-fold CV on stationary data, its capacity to preserve temporal order makes it particularly useful for many practical applications involving non-stationary data, where $k$-fold CV often falls short [38]. Therefore, in this paper, we focus on the computation of the error of the holdout estimator, deriving an analytical expression in the white inverse Wishart case. Furthermore, this result allows us to find the optimal split between the train and test sets that minimizes the Frobenius error.
@@ 2.1 Multivariate Distributions
Definition 2.2.
The sample covariance matrix $\bm{E}$ is defined as
@@
Under the assumption of normally distributed data, $\bm{E}$ is the maximum likelihood estimator of the population covariance matrix $\bm{\Sigma}$ [39].
Property 2.1.
If $\bm{x}_{1},\dots,\bm{x}_{t}$ are i.i.d. samples from $\mathcal{N}(\bm{0},\bm{\Sigma})$, then the sample covariance matrix $\bm{E}$ follows a scaled Wishart distribution:
@@
Definition 2.3.
The high-dimensional limit refers to the asymptotic regime where both $n$ and $t$ tend to infinity while the aspect ratio $q$ remains fixed
@@
Lemma 2.1.
Let $\bm{\Sigma}\sim\mathcal{W}_{n}^{-1}(t^{*},(t^{*}-n-1)\bm{\mathds{1}})$ with $t^{*}>n+1$. Then
@@
Proof.
The expected value of an inverse Wishart-distributed matrix is given by
@@
By setting $\bm{\Psi}=(t^{*}-n-1)\bm{\mathds{1}}$, we obtain $\mathbb{E}[\bm{\Sigma}]=\bm{\mathds{1}}$. ∎
@@
In this work, we will use the white inverse Wishart distribution as the model for the population covariance matrix $\bm{\Sigma}$ of multivariate normal distributed data defined in definition (2.1).
@@ 2.1.1 Estimation of Parameters
Lemma 2.2.
If $\bm{\Sigma}\sim\mathcal{W}^{-1}_{np}$ is a matrix extracted from a white inverse Wishart with parameters $\{n,p\}$ then, in the high-dimensional limit of definition  (2.3) with $p$ negligible compared to $n$, we have
@@
Proof.
The variance of the elements of $\bm{\Sigma}$ is [40]
@@
Therefore, by using the expectation of lemma (2.1), we can derive the
@@
In the high-dimensional limit, defined in definition  (2.3), the former equation simplifies to the known formula [16]
@@
Remark 2.1.
From equation  (11), $p<n/3$ is a necessary condition to have a finite variance.
To estimate $p$ from a sample covariance $\bm{E}$ with population covariance $\bm{\Sigma}$ we can use the following property.
Property 2.2.
Let $\bm{E}=\frac{1}{t}\bm{X}\bm{X}^{\top}$ be the sample covariance matrix based on $t$ i.i.d. samples from $\mathcal{N}(\bm{0},\bm{\Sigma})$, where $\bm{\Sigma}$ is the population covariance matrix. Then, in the high dimensional limit,
@@
where $q=n/t$ is the aspect ratio of the data matrix [16].
@@ 2.2.1 Oracle estimator
$\bm{\Xi}(\bm{E})$ is a RIE of the population covariance matrix $\bm{\Sigma}$, where $\bm{E}$ denote the sample covariance over the data. This estimator satisfies the following distributional equality for all orthogonal matrices $\bm{O}$[12]
[…]
@@
The Frobenius estimation error between an estimator $\bm{\Xi}$ and a population matrix $\bm{\Sigma}$ is [16]
@@
Let us mention that the expected Frobenius error of the sample covariance Wishart distributed (definition  (2.1)) is known and does not depend on the distribution of the population covariance [16].
Property 2.3.
For $\bm{E}$ a sample covariance matrix from multivariate normal data $X\in\mathbb{R}^{n\times t}$ with population covariance $\bm{\Sigma}$, in the high-dimensional limit, the expected Frobenius error between $\bm{E}$ and $\bm{\Sigma}$ is equal to
@@
In order to define the oracle estimator of the covariance matrix, let us first recall the following property.
Property 2.4.
If $\bm{E}$ is a sample covariance matrix, then it is symmetric and non-negative defined, and it can be diagonalized on an orthonormal basis. Its eigenvectors are called the sample eigenvectors and its eigenvalues are contained in the diagonal matrix $\bm{\Lambda}$ and are non-negative
@@
The RIE which minimizes the Frobenius error with the population covariance $\bm{\Sigma}$ is called the oracle estimator and is equal to [11]
@@
This estimator is called the “oracle” estimator because it still depends on the unknown population covariance matrix that we want to estimate. Ledoit and Péché derived an estimator that depends only on observable quantities and converges to the oracle in the high-dimensional limit [11].
@@
Let us call $\bm{E}$ a sample covariance of size $n\times n$ generated by $t$ i.i.d. observations with covariance population matrix $\bm{\Sigma}$. Let us assume that the noise used to generate $\bm{E}$ is independent of $\bm{\Sigma}$, has a finite $12$th absolute central moment and that the distribution of $\bm{\Sigma}$ converges for $n$ tends to $\infty$ at every point of continuity to a non-random distribution which defines a probability measure and whose support is included in a compact interval. Let $q=n/t$ and $g_{\bm{E}}$ the Stieltjes transform of $\bm{E}$. For $\lambda$, an eigenvalue of $\bm{E}$, let $\xi_{\lambda}(\bm{E})$ be the corresponding eigenvalue of the oracle’s estimator. Then, in the high-dimensional limit, for $z=\lambda+i\eta$ with $\lambda\in Sp(\bm{E})$ and $\eta>0$ [11]
@@
In reference [17], the authors claim that, even when the data matrices are finite but large, the Ledoit-Péché eigenvalues still provide a good approximation of the oracle eigenvalues.
In the high-dimensional limit for a white inverse Wishart population covariance, the Ledoit-Péché formula is a linear shrinkage that coincides with the oracle estimator. It is worth mentioning that the Bayesian estimator converges to the same formula in the high-dimensional limit [37]. In such conditions, the expected Frobenius error between the estimator and the population covariance can be derived.
Property 2.5.
For $\bm{\Sigma}$, an inverse Wishart population covariance matrix of parameters $n$ and $p$ and $\bm{E}$ the sample covariance of $\bm{\Sigma}$ generated on $t$ i.i.d. centered and Gaussian observations. Then, in the high-dimensional limit, the expected Frobenius error between the oracle estimator and $\bm{\Sigma}$ is [37, 41]
@@
An important remark is that the expected Frobenius error of the oracle estimator is smaller than the one from the sample covariance (proposition (2.3)) for any $q\neq 0$ even though the sample estimator is the maximum likelihood estimator.

### v2

@@ Abstract
Cross-validation is a statistical tool that can be used to improve large covariance matrix estimation. Although its efficiency is observed in practical applications and a convergence result towards the error of the non linear shrinkage is available in the high-dimensional regime, formal proofs that take into account the finite sample size effects are currently lacking. To carry on analytical analysis, we focus on the holdout method, a single iteration of cross-validation, rather than the traditional $k$-fold approach. We derive a closed-form expression for the expected estimation error when the population matrix follows a white inverse Wishart distribution, and we observe the optimal train-test split scales as the square root of the matrix dimension. For general population matrices, we connected the error to the variance of eigenvalues distribution, but approximations are necessary. In this framework and in the high-dimensional asymptotic regime, both the holdout and $k$-fold cross-validation methods converge to the optimal estimator when the train-test ratio scales with the square root of the matrix dimension which is coherent with the existing theory.
@@ 1 Introduction
Covariance estimation is a major question in statistics that finds applications in many areas such as physics (Taylor et al., 2013), neuroscience (Beltrachini et al., 2013; Zhuang et al., 2020), signal processing (Ottersten et al., 1998) and finance (Laloux et al., 1999; Ledoit et al., 2004; Pantaleo et al., 2011). A simple answer could be to use the sample estimator; however, it is largely inefficient in high dimensions when the order of the matrix $n$ is comparable to the number of observations $t$ of the data (Ledoit et al., 2004). More generally, it is known that the amount of information lost when the sample estimator $\bm{E}$ is used to approximate the population covariance $\bm{C}$ depends only on $q=\frac{n}{t}$ for any population matrix with Gaussian data in the high-dimension setting (Bongiorno and Lamrani, 2024). The estimator coincides with the population when $q=0$, then the larger is $q$ the larger is the error (Ledoit et al., 2004; Ledoit and Wolf, 2022a).
@@
To address the lack of robustness in the sample covariance estimator, researchers have extensively studied Rotational Invariant Estimators (RIE) (Bun et al., 2017; Ledoit and Péché, 2011; Bun et al., 2016). Today, RIEs represent the most widely used approach for covariance cleaning. In simple terms, a RIE is built by maintaining the eigenvectors of the sample covariance and correcting the sample eigenvalues. The reasoning is that when no information is available on the direction of the eigenvectors of the population covariance matrix, one can assume that each basis of eigenvectors has the same probability of appearing, therefore one keeps the sample eigenvectors and focuses on estimating the eigenvalues such that it minimizes a function of choice such as the Mean Square Error (MSE) (Ledoit and Péché, 2011), the Kullback-Leibler divergence (Bongiorno and Berritta, 2023) or the Stein’s loss (Ledoit and Wolf, 2022b). In the last decades, many RIEs linked with Random Matrix Theory (RMT) were proposed to clean the noise of the sample estimator (see (Bun et al., 2017) and (Ledoit and Wolf, 2022a) for a review). One of the first RMT methods for cleaning is known as eigenvalue clipping. This method suggests that all the eigenvalues outside the sharp limit $(\lambda_{-},\lambda_{+})$ of the Marcenko-Pastur (MP) distribution should be kept, while the ones that fall within such limit should be regarded as noise and should be shrunk to a single average value (Laloux et al., 1999; Bouchaud and Potters, 2011). However, the limits of the MP distribution depend only on the measured aspect ratio $q$, i.e., the size of the matrix $n$ divided by the number of data points $t$. Notably, the measured aspect ratio is not always equal to the effective ratio that best fits the empirical spectral density to the MP distribution (Laloux et al., 1999; Bun et al., 2017). The other widely used method for cleaning covariance matrices is linear shrinkage. The idea is to linearly shrink the sample eigenvalues towards the identity (Ledoit et al., 2004; Potters and Bouchaud, 2020a); although this approach works in many practical applications, it is theoretically justified only under some circumstances such as when the population covariance is a white inverse Wishart (Potters and Bouchaud, 2020a). More recently, the optimal shrinkage correction was analytically derived by Ledoit and Péché under certain assumptions, including the $12$th absolute central moment finite for the noise and stationary data (Ledoit and Péché, 2011); however, the computation of this quantity from their equation was not straightforward. Numerical solutions, known as Non-Linear Shrinkage (NLS), were implemented and further developed by Ledoit and Wolf (Ledoit and Wolf, 2012, 2017), making it more practical to apply. The NLS, under the high-dimensional limit, converges to the oracle estimator, which is the ideal estimator that minimizes the MSE with the population covariance matrix. It is worth remarking that the direct estimation of the oracle is unattainable in practice since it relies on the knowledge of the population covariance matrix, which is unknown (Ledoit and Péché, 2011).
@@
In practical scenarios, where the data may not be Gaussian or stationary, alternative methods inspired by machine learning find their best regime of applicability. Techniques such as Cross-Validation (CV), some based on factor models (Fama and French, 1993; Fan et al., 2008), and others employing Bayesian methods with prior distributions for estimating the covariance matrix (Daniels and Kass, 1999) can, in principle, outperform theoretically optimal methods, such as those derived by Ledoit and Péché, under these complex scenarios. However, there is no theoretical evidence to support this claim, as deriving analytical results for these methods is highly complex. To fix this gap, the focus of this work is specifically on covariance cleaning through CV.
@@
CV is one of the most widely used machine learning methods for model selection and evaluation and belongs to the family of Monte-Carlo methods (see (Berrar, 2019) for a review of the main CV procedures). Its applications include a wide range of areas such as ecology (Yates et al., 2023), medical imaging (Bradshaw et al., 2023), psychology (De Rooij and Weeda, 2020), and finance (Tan and Zohren, 2025a; Lam, 2020), to quote a few. Part of the reason CV is widely used is because it almost needs no assumptions on the data, it helps reduce the risk of over-fitting for model selection and its efficiency is undoubtedly in practice (Berrar, 2019; Tan and Zohren, 2025a). There are several CV procedures; however, the core principle of any of those involves dividing the dataset into separate training and testing subsets, also called in-sample and out-of-sample. The model is trained on the training subset and then evaluated on the testing subset. This process helps to assess how well the model generalizes to unseen data and is particularly useful to detect overfitting (Moore, 2001). When overfitting occurs, the model may perform well on the training set but will show a noticeable drop in performance on the testing set. The most popular CV methods are the $k$-fold CV, the Leave-One-Out (LOO) CV, which is a particular case of the former (Berrar, 2019), and the holdout, also called validation or simple cross-validation (Berrar, 2019). The holdout procedure consists of splitting a single time the data set into a train set and a test set. Therefore, this method consists of a single round of a $k$-fold CV. The recommendation in the literature is to use $10$ to $30\%$ of the data points for the test set and the rest for the train set (Berrar, 2019). Differently from the $k$-fold and LOO CV, the holdout can respect the causality of time series if the train and test data are ordered and if the test data is chosen posterior to the train. Practitioners should choose their approach based on the system specificity and the estimation strategy they aim for. For example, the LOO CV is adequate for bias correction but has a very high variance. On the other hand, if stability is important, one can run a $k$-fold procedure with a small k usually between $5$ and $10$, with $k=10$ being the most common one (Marcot and Hanea, 2021; Wong and Yeh, 2019; Yadav and Shukla, 2016).
@@
Covariance cleaning approaches inspired by CV were proposed in (Abadir et al., 2014; Bartz, 2016; Lam, 2016). Some authors even claim that the $k$-fold CV outperforms the NLS when applied to financial data (Bartz, 2016; Mörstedt et al., 2024). The flexibility of this numerical estimation allows for simple modifications such as including exponential smoothing in CV (Tan and Zohren, 2025a), or pre-filtering the data using hierarchical clustering (Bongiorno et al., 2022). In its original formulation, a LOO CV was proposed in Ledoit and Wolf (2012); however, Bartz (2016) and Lam (2016) reasoned that a $k$-fold would improve the estimation. The current procedure consists of dividing the dataset into $k$-folds and finding the average eigenvalue correction of the train eigenvalues to best fit the test set. Analogously to the $k$-fold problem, as $k$ increases, the train set becomes larger, but the test set shrinks, leading to a higher generalization variance. Finding the optimal way to split the data is challenging, especially because the underlying theory is not well understood. In the literature, the recommendation is again to use $k$ between $5$ and $10$ (Bartz, 2016; Bun et al., 2018; Tan and Zohren, 2025a). A first attempt to describe the CV covariance cleaning was done by Abadir et al. (2014), who studied a type of estimator, but only when the matrix dimension is fixed or becomes negligible with respect to the sample size as they tend to infinity. More recently, Lam (2016, 2015) obtained a conservative convergence result for CV estimators, showing that the Frobenius error of the holdout estimator (named NERCOME) converges to the corresponding error of the NLS. Under Lam (2016)’s assumptions, the holdout error as a function of the split should exhibit a plateau in the high-dimension regime. However, such a plateau constitutes a rough approximation in the finite-sample case, whereas our result reveals a distinct error minimum. In fact, some authors have raised concerns about the performance of these splits in the high but finite dimensional setting (Engel et al., 2017; Tan and Zohren, 2025b) and Lam (2016) himself recommends to fit the optimal split directly using the data.
@@
In this paper, we present the first step toward gaining a better understanding of the performance of CV for large but finite covariance matrix estimation using random matrix theory tools. We compute the element-wise expected MSE between the estimator and the population covariance of a single round for a $k$-fold CV, which coincides with the Frobenius error of the holdout estimator. For the general case, the error requires numerical approximations (Ledoit and Wolf, 2012, 2017; Bun et al., 2017). To carry on this estimation analytically, instead, we assume a white inverse Wishart distributed population covariances. Our motivation is to describe analytically a case with satisfies Lam’s convegence result assumptions for which we do not obtain the expected plateau for the convergence region in the high but finite dimensional regime. The inverse Wishart ensemble makes it possible to carry on analytical computations as when the datas are Gaussian, both the Bayesian estimator and linear shrinkage coincide with the oracle estimator of Ledoit and Wolf (2012) (Alvarez et al., 2014; Potters and Bouchaud, 2020b). We believe that studying the holdout method in the finite setting is a necessary step to understand the $k$-fold CV better. Furthermore, while the holdout method may be less efficient than $k$-fold CV on stationary data, its capacity to preserve temporal order could make it particularly useful for many practical applications involving non-stationary data (Bongiorno et al., 2023). Therefore, in this paper, we focus on the computation of the expected error of the holdout estimator, deriving an analytical expression in the white inverse Wishart case. Furthermore, this result allows us to find the optimal split between the train and test sets that minimizes the Frobenius error, observing that the optimal splits in the high-dimensional regime are not equivalent in the high but finite setting.
@@ 2.1 Multivariate Distributions
Then sample covariance matrix $\bm{E}$ is defined as
@@
Under the assumption of normally distributed data, $\bm{E}$ is the maximum likelihood estimator of the population covariance matrix $\bm{\Sigma}$ (Muirhead, 2009). Moreover, if $\bm{x}_{1},\dots,\bm{x}_{t}$ are i.i.d. samples from $\mathcal{N}(\bm{0},\bm{\Sigma})$, then the sample covariance matrix $\bm{E}$ follows a scaled Wishart distribution:
@@
In this paper, we will mostly consider the high-dimensional limit which refers to the asymptotic regime where both $n$ and $t$ tend to infinity while the aspect ratio $q$ remains fixed
@@
Then, for $\bm{\Sigma}\sim\mathcal{W}_{n}^{-1}(t^{*},(t^{*}-n-1)\bm{\mathds{1}})$ with $t^{*}>n+1$, we obtain that
@@
Indeed, the expected value of an inverse Wishart-distributed matrix is given by
@@
By setting $\bm{\Psi}=(t^{*}-n-1)\bm{\mathds{1}}$, we obtain $\mathbb{E}[\bm{\Sigma}]=\bm{\mathds{1}}$.
@@
In this work, we will use the white inverse Wishart distribution as the model for the population covariance matrix $\bm{\Sigma}$ of multivariate normal distributed data defined in Eq. (2).
@@ 2.1.1 Estimation of Parameters
If $\bm{\Sigma}\sim\mathcal{W}^{-1}_{np}$ is a matrix extracted from a white inverse Wishart with parameters $\{n,p\}$ then, in the high-dimensional limit of definition  (4) with $p$ negligible compared to $n$, we have
@@
To derive the normalized trace of an inverse Wishart matrix, let us use that the variance of the elements of $\bm{\Sigma}$ is (Haff, 1979)
@@
Therefore, by using the expected value of the normalized trace of $\bm{\Sigma}$ written in eq. (6), we can derive the
@@
In the high-dimensional limit, defined in definition  (4), the former equation simplifies to the known formula (Potters and Bouchaud, 2020a)
@@
Let us note that from equation  (11), $p<n/3$ is a necessary condition to have a finite variance.
To estimate $p$ from a sample covariance $\bm{E}$ with population covariance $\bm{\Sigma}$ we can use the expected value of the trace of $\bm{E}^{2}$. Let $\bm{E}=\frac{1}{t}\bm{X}\bm{X}^{\top}$ be the sample covariance matrix based on $t$ i.i.d. samples from $\mathcal{N}(\bm{0},\bm{\Sigma})$, where $\bm{\Sigma}$ is the population covariance matrix. Then, in the high dimensional limit,
@@
where $q=n/t$ is the aspect ratio of the data matrix (Potters and Bouchaud, 2020a).
@@ 2.2.1 Oracle estimator
$\bm{\Xi}(\bm{E})$ is a RIE of the population covariance matrix $\bm{\Sigma}$, where $\bm{E}$ denote the sample covariance over the data. This estimator satisfies the following distributional equality for all orthogonal matrices $\bm{O}$(Bun et al., 2016)
$\bm{\Xi}(\bm{O}\bm{E}\bm{O}^{T})\overset{\textrm{distr}}{=}\bm{O}\Xi(\bm{E})\bm{O}^{T}.$ (15)
@@
The normalized Frobenius estimation error between an estimator $\bm{\Xi}$ and a population matrix $\bm{\Sigma}$ is (Potters and Bouchaud, 2020a)
@@
Let us mention that the expected Frobenius error of the sample covariance Wishart distributed is known and does not depend on the distribution of the population covariance (Potters and Bouchaud, 2020a).For $\bm{E}$ a sample covariance matrix from multivariate normal data $X\in\mathbb{R}^{n\times t}$ with population covariance $\bm{\Sigma}$, in the high-dimensional limit, the expected Frobenius error between $\bm{E}$ and $\bm{\Sigma}$ is equal to
@@
In order to define the oracle estimator of the covariance matrix, let us first recall that $\bm{E}$ can be diagonalized on an orthogonal basis. Its eigenvectors are called the sample eigenvectors and its eigenvalues are contained in the diagonal matrix $\bm{\Lambda}$ and are non-negative
@@
The RIE which minimizes the Frobenius error with the population covariance $\bm{\Sigma}$ is called the oracle estimator and is equal to (Ledoit and Péché, 2011)
@@
This estimator is called the “oracle” estimator because it still depends on the unknown population covariance matrix that we want to estimate. Ledoit and Péché derived an estimator that depends only on observable quantities and converges to the oracle in the high-dimensional limit (Ledoit and Péché, 2011).
@@
Let us call $\bm{E}$ a sample covariance of size $n\times n$ generated by $t$ i.i.d. observations with covariance population matrix $\bm{\Sigma}$. Let us assume that the noise used to generate $\bm{E}$ is independent of $\bm{\Sigma}$, has a finite $12$th absolute central moment and that the distribution of $\bm{\Sigma}$ converges for $n$ tends to $\infty$ at every point of continuity to a non-random distribution which defines a probability measure and whose support is included in a compact interval. Let $q=n/t$ and $g_{\bm{E}}$ the Stieltjes transform of $\bm{E}$. For $\lambda$, an eigenvalue of $\bm{E}$, let $\xi_{\lambda}(\bm{E})$ be the corresponding eigenvalue of the oracle’s estimator. Then, in the high-dimensional limit, for $z=\lambda+i\eta$ with $\lambda\in Sp(\bm{E})$ and $\eta>0$ (Ledoit and Péché, 2011)
@@
Ledoit and Wolf (2012) claim that, even when the data matrices are finite but large, the Ledoit-Péché eigenvalues still provide a good approximation of the oracle eigenvalues.
In the high-dimensional limit for a white inverse Wishart population covariance, the Ledoit-Péché formula is a linear shrinkage that coincides with the oracle estimator. It is worth mentioning that the Bayesian estimator converges to the same formula in the high-dimensional limit (Potters and Bouchaud, 2020b). In such conditions, the expected Frobenius error between the estimator and the population covariance can be derived.
Property 2.1.
For $\bm{\Sigma}$, an inverse Wishart population covariance matrix of parameters $n$ and $p$ and $\bm{E}$ the sample covariance of $\bm{\Sigma}$ generated on $t$ i.i.d. centered and Gaussian observations. Then, in the high-dimensional limit, the expected Frobenius error between the oracle estimator and $\bm{\Sigma}$ is (Potters and Bouchaud, 2020b; Lamrani, 2021)
@@
An important remark is that the expected Frobenius error of the oracle estimator is smaller than the one from the sample covariance written in equation (17) for any $q\neq 0$ even though the sample estimator is the maximum likelihood estimator.